OpenStudy (anonymous):

sin theta = cos theta/ root 3 I need theta but the inverse of this is a maths error...

OpenStudy (anonymous):

\[\sin \theta = \frac{ \cos \theta }{ \sqrt{3} } ?\]

OpenStudy (anonymous):

sorry why are there all the lines I find it hard to read

OpenStudy (anonymous):

@antoni7

OpenStudy (anonymous):